karissafrazier:

what's 6x+10x divided by 2x+9x?

jhonyy9:

there are sure 6x+10x and 2x+9x ? any x without any exponent ?

karissafrazier:

look, idk, i'm trying to get smarter with my geometry

jhonyy9:

geometry ??? but this is not geometry

karissafrazier:

geometry is mathmatics

jhonyy9:

ok but this your problem is algebra math so 6x+10x = ? there are the same terms what you can assume them same with 2x+9x = ?

karissafrazier:

16x divided by 11x

jhonyy9:

yes of course

jhonyy9:

\[\frac{ 16x }{ 11x } = ?\]

jhonyy9:

what are the same terms in numerator and denominator ? what can you make with these terms of x ?

karissafrazier:

got this

jhonyy9:

how you get this result ?

jhonyy9:

\[\frac{ 16x }{ 11x } = \frac{ 16 }{ 11 }\]

jhonyy9:

bc. you can simplifie the x from the top and down

jhonyy9:

ok ? any questions ?

karissafrazier:

no

karissafrazier:

i do want a reply lock on this tho

jhonyy9:

there is the reply ...